OpenStudy (anonymous):

Find the vertex of the parabola for the equation: 3x2 - 6x - 5 a.(1, -4) b.(1, -8) c.(-1, -4) d.(-1, -8)

OpenStudy (nenadmatematika):

do the first derivative and set it to be zero, and you'll get y'=6x-6=0 solve it for x; x=1and then plug it back into the equation and you'll get y: y=3-6-5=-8 so your vertex is (1,-8)